OpenStudy (anonymous):

A block and tackle is used to raise a 72.0 N crate 12.3 m from the loading dock to the ship. This requires a force of 25.0 N to be applied by the longshoreman. Part 1: What is the mechanical advantage of the block and tackle? Part 2: How much rope must be pulled? Note: Assume that IMA = MA. Part 3: The longshoremen have only a 40 m rope that they are told not to cut. They use this for the block and tackle. How does this affect the efficiency of the block and tackle? **how do we solve these? :/ thanks @Michele_Laino !!:)

OpenStudy (michele_laino):

part 1 the requested advantage is given by the subsequent ratio: \[\large R = \frac{{72}}{{25}}\]

OpenStudy (anonymous):

oh okay! so we get 2.88 is the requested mechanical advantage?

OpenStudy (michele_laino):

yes! Whereas the efficiency E, of our machine, is: \[\Large E = \frac{{25}}{{72}} = ...\]

OpenStudy (anonymous):

ohh okay! it is 0.347222....

OpenStudy (michele_laino):

no, it is an additional information only. Our answer, to part 1 is 2.88

OpenStudy (anonymous):

oh ok! so we do not need to include the efficiency? just the 2.88 mechanical advantage?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

okie! so onto part 2!

OpenStudy (michele_laino):

for part #2, if I call with L the length of the pulled rope, then we can write: \[\large 72 \times 12.3 = 25 \times L\]

OpenStudy (michele_laino):

so we have: \[\Large L = \frac{{72 \times 12.3}}{{25}} = ...meters\]

OpenStudy (anonymous):

we get 35.424!

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

so is that the amount of how much rope must be pulled? 35.424 m?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

ok! yay!! so onto part 3?

OpenStudy (michele_laino):

if we have to pull 40 meters of rope, then the force F applied by the longshoremen, is: \[\Large F = \frac{{72 \times 12.3}}{{40}} = ...Newtons\]

OpenStudy (anonymous):

ok! so we get 22.14 Newtons! is that our solution for part 3?

OpenStudy (anonymous):

oh wait, that is just the force, right? how can we find the efficiency?

OpenStudy (michele_laino):

the new advantage is: \[\Large R = \frac{{72}}{{22.14}} = ...\]

OpenStudy (anonymous):

or how can we find how it affects the efficiency?

OpenStudy (michele_laino):

whereas the new efficiency, is: \[\Large E = \frac{{22.14}}{{72}} = ...\]

OpenStudy (michele_laino):

I think that your problem asks for efficiency, also in part #1

OpenStudy (anonymous):

ok! so we get 0.3075 is the efficiency? and i think we don't need it for part 1? since it asks for mechanical advantage only? or do we need to include the efficiency we calculated earlier for part 1?

OpenStudy (michele_laino):

so the answer to part #1 is: \[\Large E = \frac{{25}}{{72}} \cong 0.35\]

OpenStudy (michele_laino):

that's right! the new efficiency is: \[\Large E = \frac{{22.14}}{{72}} \cong 0.31\]

OpenStudy (michele_laino):

namely the efficiency will decrease

OpenStudy (anonymous):

yes:) and yay!! so that is how it affects efficiency? it will decrease?

OpenStudy (michele_laino):

yes! correct!

OpenStudy (anonymous):

yay! thanks!! so we are done?

OpenStudy (michele_laino):

yes! we have finished!

OpenStudy (anonymous):

yay!! ok! onto the next!!:)

OpenStudy (anonymous):

thank you!!

OpenStudy (michele_laino):

:) :)